OpenStudy (anonymous):

i am stumped on this one....solve w^4-18w^2-2=0

OpenStudy (amistre64):

rename the variables; w^2 = x x^2 -18x -2 = 0

OpenStudy (amistre64):

(x^2 -18x +81) -2 = 81 ; completed a square (x-9)^2 = 83 ; compacted it all x-9 = +- sqrt(83) ; cant remember what i did here x = 9 +- sqrt(83)

OpenStudy (amistre64):

now; x = w^2 sooo; w^2 = 9+- sqrt(83) ; solve for w